OpenStudy (anonymous):

[7.03] Compare and contrast, in complete sentences, the difference between the products of a^2a^4 and (a^2)^4. Provide the expanded form and product of each.

OpenStudy (anonymous):

a^2a^4 = a^(2+4)=a^6 (a^2)^4=a^(2*4)=a^8

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

ok but whats similar about them

OpenStudy (chihiroasleaf):

\[a ^{2}timesa ^{4}=a ^{2+4}=a ^{6}\] \[\left( a ^{2} \right)^{4}=a ^{2\times4}=a ^{8}\]

OpenStudy (anonymous):

whats similar

OpenStudy (anonymous):

???

OpenStudy (anonymous):

They have no similarity. i.e they are NOT the same. except, maybe, that a^8=(a^6)*(a^2)?